Consider the following information for stocks A, B, and C. The returns on the three stocks are positively correlated, but they are not perfectly correlated. (That is, each of the correlation coefficients is between 0 and 1.)

Stock Expected Return Standard Deviation Beta
A 8.19% 14% 0.7
B 10.46    14    1.2
C 11.83    14    1.5

Fund P has one-third of its funds invested in each of the three stocks. The risk-free rate is 5%, and the market is in equilibrium. (That is, required returns equal expected returns.)

  1. What is the market risk premium (rM - rRF)? Round your answer to two decimal places.
      %
  2. What is the beta of Fund P? Do not round intermediate calculations. Round your answer to two decimal places.

  3. What is the required return of Fund P? Do not round intermediate calculations. Round your answer to two decimal places.
      %
  4. Would you expect the standard deviation of Fund P to be less than 14%, equal to 14%, or greater than 14%?
    1. Less than 14%
    2. Greater than 14%
    3. Equal to 14%
Answer #1

As per CAPM Model,

Required Rate = Rf + Beta(Market Risk Premium)

Market Risk Premium = (0.0819 - 0.05)/0.7

Market Risk Premium =4.56%

Beta of Fund P = 0.33(0.70) + 0.33(1.2) + 0.33(1.5)

Beta of Fund P = 1.13

Required Rate of Fund P = 0.05 + 1.13(0.0456)

Required Rate of Fund P = 10.15%

Standard Deviation of Fund P will be greater than 14% as there will be element of correlation between stocks in the fund, which will contribute to standard deviation along with standard deviation of individual stocks.